# Analyzing Logs of Large-Scale Software Systems using Time Curves Visualization

### v1

@@ III-G Distance computation
We include the normalization by weights to further generalize the proposed distance, allowing a custom weighting to favor certain types of editions if prior information about the logging system exists. However, by default the proposed method uses $w_{ins}=w_{del}=w_{sub}=1$ , meaning the normalization by weights is not necessary.
The next step is aggregation. Our proposal consists of iterating though all templates of one checkpoint to find the closest template in the other checkpoint and collect the distances.
To maintain 0-1 scale, we employ the arithmetic mean of all the collected distances between templates of checkpoint $x$ and $y$:
@@
We iterate and average over templates of the first checkpoint (and not the second), meaning $D^{\prime}$ is not symmetric. A simple solution is $D(x,y):=D(y,x):=\frac{D^{\prime}(x,y)+D^{\prime}(y,x)}{2}$. However, such a measure disregards cardinalities of checkpoints, making it too sensitive to checkpoints with few templates and thus not robust. After multiple experiments, a following logarithmic weighting was chosen for the final distance function:
@@
Note that denominator is used just for normalization, i.e.:
$D^{\prime}(x,y),D^{\prime}(x,y)\in[0,1]\implies D(x,y)=D(y,x)\in[0,1].$
The proposed distance function is formally not a metric, since a metric needs to satisfy the following properties for any elements $x$ and $y$ from the set:
  1. 1.
    Non-negativity: $dist(x,y)\geq 0$.
  2. 2.
    Identity of indiscernibles: $dist(x,y)=0\iff x=y$.
  3. 3.
    Symmetry: $dist(x,y)=dist(y,x)$.
  4. 4.
    Triangle inequality: $dist(x,z)\leq dist(x,y)+dist(y,z)$.
If only the first 3 properties are satisfied, but triangular inequality is violated, then $dist$ is a semimetric. We will now prove that $D(\cdot,\cdot)$ as defined in Eq. 2 is indeed a semimetric over the set of all possible checkpoints.
  1. 1.
    Since all terms are non-negative, so is the $D(\cdot,\cdot)$, i.e.:
    $D^{\prime}(x,y)=\frac{1}{m}\sum_{i=0}^{m}\min_{j}d(y_{i},x_{j})\overset{d(x_{i},y_{j})\geq 0}{\geq}0.$ (3)
[…]
  2. 2.
    Identity of indiscernibles can be verified as follows:
    $D(x,y)=0$ $\iff D^{\prime}(x,y)=D^{\prime}(y,x)=0$
    $\iff\left\{\begin{array}[]{ll}\forall i\min_{j=0,\dots,n-1}&d(x_{i},y_{j})=0,\\ \forall j\min_{i=0,\dots,m-1}&d(y_{j},x_{i})=0.\\ \end{array}\right.$
    $\iff\forall i\ \ x_{i}\in y\text{ and }\forall j\ \ y_{j}\in x$
    $\iff x\subseteq y\text{ and }y\subseteq x$
    $\iff x=y.$
  3. 3.
    Symmetry holds by design, because $D(x,y)$ involves both $D^{\prime}(x,y)$ and $D^{\prime}(y,x)$.
  4. 4.
    The triangular inequality is the only violation of metric properties. This can be seen with nested checkpoints (a counterexample is available in the Supplementary Material [22]).
Overall, the proposed semimetric requires significant computation since it comprises multiple quadratic terms. Further details on computational optimizations are available in [22].
III-H Changes to the original Time Curves
@@
From the user interface, we replaced the color scale in order to account for time. In addition, we included an animation option to draw the Time Curve segment-by-segment from start to end. This helps to understand the evolution of the system by slowly incorporating events, especially as projections get more complex. Example animations can be found in [22]. Lastly, we included the coefficient of determination ($R^{2}$) as an indicator of the projection quality.
From the data augmentation perspective, we enriched data points by including the log template messages contained within as annotations that can be opened and closed via clicking. This enables the user to explore and compare the log messages of different instants of the system. Furthermore, this facilitates post-processing via e.g., LLMs.
From the functionality perspective, we included the possibility of overlaying multiple curves together, which can be used to compare multiple similar systems at once (see e.g., [23]). More details can be found in the results and discussion in Section IV-D.
@@ IV Evaluation
Figure 2: Projections of the collected logs for three stream processing framework executions and their unfolding according to the weight of the time dimension. The time influence is the proportion of the linear interpolation between similarity and time. The annotated labels, from A to D represent states of interest for the analysis of the system, namely (A) startup, (B1-B5) failure injections, (C1-C5) recovery and (D) shutdown. Two additional cases are also annotated, normal processing after recovery as labels C1’-C5’ and a singular event in Kafka Streams marked with $\ast$.
In order to evaluate the method, we conducted multiple experiments, focusing on explainability and scalability. First, to confirm the expected behavior of the proposed approach, a controlled experiment was repeated on three different stream processing frameworks and the output logs were analyzed. The objective is to achieve a consistent analysis across all frameworks despite their logging differences. We also analyzed a publicly-available collection of logs first manually and then leveraged the summarizing capabilities of LLMs for single and pairwise checkpoints. This is aimed at showing how the visual interpretation of the Time Curves can be used to gather insights while also enriching with context information. In addition, we analyzed a proprietary log dataset of three concurrent systems to showcase the multi-curve visualization. Lastly, for the case of scalability, we measured the performance of the method for a variety of publicly-available log datasets of different sizes.
@@ IV-A Stream processing analysis
While many public log datasets are available in the literature, they are hardly annotated or explained, thus making it very difficult to analyze without bias. Therefore, we reused the log data of our previous research on fault recovery [24]. This provides a ground-truth collection of logs from different applications executing the same task, enabling a fair comparison of the logs. This includes three popular stream processing frameworks, namely Apache Flink [25], Spark Structured Streaming [26] and Kafka Streams [27]. The expectation is that the proposed method must result in the same analysis outcome for all cases.
@@
The logs include a 72 minute execution of the ShuffleBench benchmark [28] where 5 failures were automatically injected every 12 minutes, starting at the 12-minute mark. Default parameters of the event detection algorithm were used except for Kafka Streams, which required to fine-tune the time granularity due to the fact that the collector node received nearly no logs from the failing nodes. Further details on the experimental setup are available on [24, 22].
@@
Figure 2 shows the Time Curves projections of each of the executions of the frameworks in three different states, either with a low, medium or high importance of the time dimension. This is provided because of the limitations of showing interactive plots within a static manuscript. Animations can be found in [22]. In addition, Fig. 2 also includes the labeling of the processing stages that take place throughout the execution, namely (A) startup, (B1-B5) failure injections, (C1-C5) fault recovery and (D) shutdown of the system.
@@
Besides the starting and shutdown sequence, notice how the upper row of Section IV, which corresponds to the exclusively similarity-based projection, shows a cyclic behavior. All three frameworks exhibit a linearly separable structure that splits failures from recoveries or normal processing, which is shown with a manually-added dashed-black line. This becomes more apparent as the curve gets unfolded by time in the middle and lower rows. Due to the proximity between the labels of failure and recovery in Flink, the normal processing state is also highlighted through labels (C1’-C5’).
@@
Example logs for each of the states can be found in [22]. Both the startup (A) and shutdown (D) can be easily identified since they contain unique messages. This is the case for all frameworks except for Kafka Streams, where there is no actual record corresponding to the shutdown sequence. In fact, label D contains only generic information messages regarding normal processing, which is no different from the previous nodes that take place after recovery in C5. This is due to how the execution was conducted and the architecture of Kafka Streams. In particular, the lack of a manager node and how the nodes are terminated after the benchmark execution result in no specific logs for the shutdown.
@@
Similarly, both Spark and Flink curves transition much faster from startup to the first injected failure (A to B1), with Spark including a single checkpoint in between and Flink four additional ones. Note that the word checkpoint is used as in a data point in the figure, rather than the checkpointing process in stream processing frameworks. On the contrary, Kafka Streams is separated into two different patterns: an initial loop right after startup and a cyclic loop afterwards. The first one is due to the initial partition assignments and balancing happening before stable processing is achieved. The second one is due to the recurrent failure injections. Interestingly, it can be seen that the projection places the cluster of failures (labels B1-B5 close to the tightly coupled checkpoints of the initial partition assignment (marked with $\ast$). This is due to the large overlap in terms of logs between the initial partition assignment and the failure injections, since after a failure a re-balancing and new partition assignment is triggered [24].
@@
A higher influence of time (i.e., bottom row) provides with a timeline representation of the sequence of events. In fact, the periodicity of the events (every 12 minutes) can be seen in terms of constant distance between the failures. The time to recovery can also be inferred, showing that in most cases it also remains quite constant, although different for each framework. Notice that further analysis is possible by increasing or decreasing the checkpoint granularity.
These results confirm our expectations based on our previous research [24], particularly by capturing the cyclic structure of failures and recoveries. It also provides additional information such as the mirrored loop of the Kafka Streams execution.
@@ IV-B Public dataset analysis
Figure 3 shows the Time Curve for the dataset along with identified sections of interest, including labels A to F and the left, middle and right plane sections. These will be used to describe the events taking place throughout the logs.
In general, the Zookeeper service begins with startup, where multiple rounds of elections take place to choose the leader server. These include logs regarding communication between the three existing servers. Once a server is elected as a leader, clients may start connecting. Such connections may be successful or not and, in addition, servers may crash due to unknown circumstances and therefore trigger new elections. The process continues for roughly 26 days according to the logs collected.
@@
    Left section includes startup sequence (A), initial configuration attempts (B) and end sequence (F), but does not contain any client requests and depicts that the system is not in a ready state.
@@
    Middle section shows relative stability periods in which the system is servicing client requests, however it is intertwined with election procedures either routinely (C) or due to occasional crashing of a server (E).
@@
    Right section depicts high stability periods in which no election procedures are required, clients are connecting and requests are being fulfilled (D). No server crashes take place in these states.
In addition, a cyclic pattern between the relative and high stability periods (C and D) can be observed, both in terms of events and time (see drastic change in color between data points in D, which denotes a large time gap).
@@
    Label A corresponds to the startup of the Zookeeper service. This involves log records such as the main thread starting the application and reading configuration files. Note that label F (and its surrounding data points) are close to A due to lack of client connections. In addition, the error “Cannot open channel to ... at election address” prevents the election to complete successfully, thus triggering further elections processes and therefore resembling the startup sequence. Such error gets resolved and the election completes at F.
@@
    Label B shows an interesting case: after startup and initial election, the system enters a repetitive state where servers are trying to communicate with each other but are unsuccessful. An unspecified error results in broken connections for all servers, i.e., “Connection broken for id ... error = ”. This behavior takes place for roughly half an hour and is responsible for almost 75% of all the log records in the dataset. Note that despite the over representation of B, the proposed method is still able to capture several other events and behaviors.
@@
    Label D shows the period of the system where highest stability is achieved without server crashes nor election procedures. Only client connections take place. Again, the variability is explained by small differences such as occasional exceptions that do not result in crashes.
@@
    Label E shows a large time gap between D of nearly 8 days of limited activity. The connection of clients resumes with occasional crashing prior to returning back to the stable section in D. The main difference between E and the cluster of data points at C is that new errors appear regarding the Zookeeper service not running temporarily.
Figure 3: Identified sections of interest corresponding to the Zookeeper log dataset. Labels A to F represent points of interest. Two planes marked with a gradient color separate the projection into left, middle and right sections.
@@ IV-C Enrinchment via LLM
When the collection of logs comprises a large number of templates (e.g., hundreds of templates in a single checkpoint), manual inspection as in Section IV-B becomes difficult due to the amount of information contained. As each point is represented by string templates, given context and domain source, LLMs can be used to further summarize the content of a single checkpoint as well as compare different data points of the Time Curve. For the following analysis, we employed Microsoft’s Copilot model which is based on GPT-4 [29].
@@ IV-C1 Single point summarization
The objective is to explain the templates contained in a single checkpoint. This can be helpful for a variety of reasons, e.g., adding context to the meaning of log records, especially when users are not familiar with the semantics of the logging format, further summarizing a large number of templates into few sentences or reasoning and inquiring about particular log messages, etc.
Two arbitrary checkpoints are summarized from labels C and D. In label C, which corresponds to a period of instability, the model explained that “The logs indicate that the ZooKeeper server is actively managing client connections and sessions, handling leader elections, initializing and configuring the server, and synchronizing data with followers. The server also deals with various warnings and errors, such as unexpected shutdowns, broken connections, and session expirations”. Note that this captures all of the manual explanation given in the previous section. When further inquired about the gravity of the errors, the model indicates that they should be investigated due to their severity.
On the contrary, in label D (stability period), the model explained described that “The logs indicate that the ZooKeeper server is actively managing client connections and sessions. It handles new connection attempts, establishes and terminates sessions, and processes session expirations. The server also deals with connection issues, such as end-of-stream exceptions and connection requests from outdated clients.”. This is in line with our analysis on the previous section. Note that the end-of-stream exceptions do not result in server crashes. Further inquiring about the seriousness of the end-of-stream exceptions result in the model explaining that they are relatively common and not problematic, as opposed to the checkpoint in Label C.
@@ IV-C2 Pairwise point comparison
The objective of pairwise comparison is to describe the differences (or similarities) between two given checkpoints. This is helpful when comparing the state of the system at two time frames or to correlate visual differences in the Time Curve with the actual set of logs.
First, we selected two close checkpoints from Label D (stable processing period) to find out if the difference is significative. The model explains that both checkpoints are nearly identical except for the presence of a single additional log record related to a KeeperException where an attempt to create an already existing node was made. When further inquiring, the model describes the exception as a non-critical error which could be handled more gracefully. This is in agreement with our manual investigation, where we found out that no server crashes occur in the high-stability period of Label D.
Second, we selected Label C and E from the unstable procession period. On one hand, regarding the similarities, the model explains that both are “… capturing various operational events and issues related to leader election, session management, and client connections”. On the contrary, regarding the differences, first the model returns a generic (yet true) summarization, namely “the first set of logs includes more detailed information about the server environment and configuration, while the second set includes more details about the election process and issues encountered during synchronization and session management.” along with a list of the differences depending on their severity level (i.e., variations by INFO, WARN, etc.). When requested that a more specific explanation is given for the set of WARN logs, the model replies “These logs highlight issues related to leader election and synchronization, such as unexpected zxid values and exceptions when following the leader, as well as a specific session closure due to the server not running”, which is in line with our manual analysis where we found out that the Zookeeper service is not running temporarily.

### v2

@@ III-F Distance computation
We include the normalization by weights to further generalize the proposed distance, allowing a custom weighting to favor certain types of editions if prior information about the logging system exists. However, since weights are one by default, no weight normalization is required.
The next step is aggregation. Our proposal consists of iterating through all templates of one checkpoint to find the closest template in the other checkpoint and collect the distances.
The arithmetic mean of all the collected distances between templates of checkpoint $x$ and $y$ is used to stay within the range $[0,1]$:
@@
For $D^{\prime}(x,y)$ we iterate and average over templates of $x$ whereas for $D^{\prime}(y,x)$ – over $y$, therefore $D^{\prime}$ is not symmetric. However, solving with $D(x,y):=D(y,x):=\frac{D^{\prime}(x,y)+D^{\prime}(y,x)}{2}$ disregards cardinalities of checkpoints, making the measure too sensitive to checkpoints with few templates. In order to make the distance more robust to the number of templates, the following logarithmic weighting was chosen:
@@
The proposed distance function is a semimetric (and not a metric) since the triangular inequality is violated, however, all other properties hold [8]. Note that this semimetric requires significant computation since it comprises multiple quadratic terms, so actual calculation differs from definition. Further details on computational optimizations are available in [8].
III-G Changes to the original Time Curves
@@
From the user interface, we replaced the color scale in order to account for time. In addition, we included an animation option to draw the Time Curve segment-by-segment from start to end. This helps to understand the evolution of the system by slowly adding checkpoints, especially as projections get more complex. Example animations can be found in [8]. Lastly, we included the coefficient of determination ($R^{2}$) as an indicator of the projection quality.
From the data augmentation perspective, the data points are enriched by including the log template messages contained within as annotations that can be opened and closed via clicking. This enables the user to explore and compare the log messages of different instants of the system. Furthermore, this facilitates further analysis via e.g., LLMs.
From the functionality perspective, we included the possibility of overlaying multiple curves together, which can be used to compare multiple similar systems at once (e.g., [23]). More details can be found in Section IV-D.
@@ IV Evaluation
Figure 3: Projections of the collected logs for three stream processing framework executions and their unfolding according to the ratio of similarity and time influences. The annotated labels (A to D) represent checkpoints of interest for the analysis of the system, namely (A) startup, (B1-B5) failure injections, (C1-C5) recovery and (D) shutdown. In addition, C1’-C5’ mark normal processing after recovery and a $\ast$ represents a singular checkpoint in Kafka Streams.
In order to evaluate the method, we conducted multiple experiments, focusing on explainability and scalability. First, to confirm the expected behavior of the proposed approach, an experiment was repeated under the same conditions on three different stream processing frameworks, and the output logs were analyzed. The objective is to achieve a consistent analysis across all frameworks despite their logging differences. We also analyzed a publicly-available collection of logs first manually and then leveraged the summarizing capabilities of LLMs for single and pairwise checkpoints. This is aimed at showing how the visual interpretation of the Time Curves can be used to gather insights while also enriching with context information. In addition, we analyzed a proprietary log dataset of three concurrent systems to showcase the multi-curve visualization. Lastly, for the case of scalability, we measured the performance of the method for a variety of publicly-available log datasets of different sizes.
@@ IV-A Stream processing analysis
While many public log datasets are available in the literature, they are hardly annotated or explained, thus making it very difficult to analyze without bias. Therefore, we reused the log data of our previous research on fault recovery [24], which includes three popular open-source stream processing frameworks, namely Apache Flink [25], Spark Structured Streaming [26] and Kafka Streams [27]. This provides an explainable collection of logs from different applications executing the same workload, enabling a comparison of the logs. The expectation is that the proposed method results in a similar analysis outcome for all cases.
@@
The logs include a 72 minute execution of the ShuffleBench benchmark [28] where 5 failures were automatically injected every 12 minutes, starting at the 12-minute mark. Default parameters of the event detection algorithm were used, except for the case of Kafka Streams, which required to fine-tune the time granularity due to the fact that the collector node received nearly no logs from the failing nodes. Further details on the experimental setup are available in [24, 8].
@@
Figure 3 shows the Time Curves projections of each of the executions of the frameworks in three different ways: with a low, medium and high importance of the time dimension. This is provided because of the limitations of showing interactive plots within a static manuscript. Animations can be found in [8]. In addition, Fig. 3 also includes the labeling of the processing stages that take place throughout the execution, namely (A) startup, (B1-B5) failure injections, (C1-C5) fault recovery and (D) shutdown of the system.
@@
Besides the starting and shutdown sequences, notice how the upper row of Fig. 3 (which corresponds to the projection based exclusively on similarity) shows a cyclic behavior. All three frameworks exhibit a linearly separable structure that splits failures from recoveries or normal processing, which is shown with a manually-added dashed-black line. This becomes more apparent as the curve gets unfolded by time in the middle and lower rows. Due to the proximity between the labels of failure and recovery in Flink, the normal processing checkpoints are also highlighted through labels (C1’-C5’).
@@
Example logs for each of the checkpoints can be found in [8]. Both the startup (A) and shutdown (D) can be easily identified since they contain unique messages. This is the case for all frameworks except for Kafka Streams, where there is no actual record corresponding to the shutdown sequence. In fact, D contains only generic information messages regarding normal processing, which is no different from the previous checkpoints that take place after recovery in C5. This is due to how the execution was conducted and the architecture of Kafka Streams. In particular, the lack of a manager node and how the nodes are terminated after the benchmark execution result in no specific logs for the shutdown.
@@
Similarly, both Spark and Flink curves transition much faster from startup to the first injected failure (A to B1), with Spark including a single checkpoint in between and Flink four additional ones. Note that the word checkpoint is used as a data point in the figure, rather than the checkpointing process in stream processing frameworks. On the contrary, Kafka Streams is separated into two different patterns: an initial loop right after startup and the main loop afterwards. The first one is due to the initial partition assignments and balancing happening before stable processing is achieved. The second one is due to the recurrent failure injections. Interestingly, it can be seen that the projection places the cluster of failures (labels B1-B5 close to the tightly coupled checkpoints of the initial partition assignment (marked with $\ast$). This is due to the large overlap in terms of logs between the initial partition assignment and the failure injections, since after a failure a re-balancing and new partition assignment is triggered [24].
@@
A higher influence of time (i.e., bottom row) provides a more linear representation of the sequence of events. In fact, the periodicity of the system (every 12 minutes) can be seen in terms of constant distance between the failures. The time to recovery can also be inferred, showing that in most cases it also remains quite constant, although different for each framework. Notice that further analysis is possible by increasing or decreasing the checkpoint granularity.
These results confirm our expectations based on our previous research [24], particularly by capturing the cyclic structure of failures and recoveries. It also provides additional information such as the unexpected startup loop of the Kafka Streams execution.
@@ IV-B Public dataset analysis
Figure 4 shows the Time Curve for the dataset along with identified sections of interest, including labels A to F and the left, middle and right plane sections. These will be used to describe the processes taking place throughout the logs.
In general, the Zookeeper service begins with startup, where multiple rounds of elections take place to choose the leader server. These produce logs about communication between the three existing servers. Once a leader is elected, clients may start connecting with or without success. In addition, servers may crash due to unknown circumstances and therefore trigger new elections. This process continued for roughly 26 days according to the collected logs.
@@
    Left section includes startup (A), initial configuration attempts (B) and end (F), but contains no client requests and depicts that the system is not in a ready state.
@@
    Middle section shows relative stability stages in which the system is servicing client requests, however it is intertwined with election procedures either routinely (C) or due to occasional crashing of a server (E).
@@
    Right section depicts high stability stages in which no election procedures are required, clients are connecting and requests are being fulfilled (D). No server crashes take place during these checkpoints.
In addition, a cyclic pattern between the relative and high stability stages (C and D) can be observed, both in terms of events and time (see drastic change in color between data points in D, which denotes a large time gap).
@@
    Label A corresponds to the startup of the Zookeeper service. This involves log records such as the main thread starting the application and reading configuration files. Note that F (and its surrounding data points) are close to A due to lack of client connections. In addition, the error "Cannot open channel to ... at election address" prevents the election to complete successfully, thus triggering further elections processes and therefore resembling the startup sequence. Such error gets resolved and the election completes at F.
@@
    Label B shows an interesting case: after startup and initial election, the system enters a repetitive phase where servers are trying to communicate with each other but are unsuccessful. An unspecified error results in broken connections for all servers, i.e., "Connection broken for id ... error = ". This behavior takes place for roughly half an hour and is responsible for almost 75% of all the log records in the dataset. Note that despite the overrepresentation of B, the proposed method is still able to capture several other behaviors.
@@
    Label D shows the stage of the system where highest stability is achieved without server crashes or election procedures. Only client connections take place. Again, the variability is explained by small differences such as occasional exceptions that do not result in crashes.
@@
    Label E shows a large time gap up until D of nearly 8 days of limited activity. The connection of clients resumes with occasional crashing prior to returning back to the stable stage in D. The main difference between E and the cluster of data points at C is that new errors appear regarding the Zookeeper service not running temporarily.
Figure 4: Identified sections of interest corresponding to the Zookeeper log dataset. Labels A to F represent points of interest. Two planes marked with a gradient color separate the projection into left, middle and right sections, corresponding to different stability stages of the system.
@@ IV-C Enrinchment via LLM
When the collection of logs comprises a large number of templates (e.g., hundreds of templates in a single checkpoint), manual inspection as in Section IV-B becomes difficult due to the amount of information contained. As each point is represented by string templates, given context and domain source, LLMs can be used to further summarize the content of a single checkpoint as well as compare different data points of the Time Curve. For the following analysis, we used Microsoft’s Copilot model which is based on GPT-4 [29].
@@ IV-C1 Single point summarization
The objective is to explain the templates contained in a single checkpoint. This can be helpful for a variety of reasons, e.g., adding context to the meaning of log records, especially when users are not familiar with the semantics of the logging format, further summarizing a large number of templates into few sentences or reasoning about particular log messages, etc.
Two arbitrary checkpoints are summarized from stages C and D. The model explained stage C as “The logs indicate that the ZooKeeper server is actively managing client connections and sessions, handling leader elections, initializing and configuring the server, and synchronizing data with followers. The server also deals with various warnings and errors, such as unexpected shutdowns, broken connections, and session expirations”. Note that this captures our previous manual explanations. When further inquired about the gravity of the errors, the model indicates that they should be investigated due to their severity, and that they are not part of what can be considered normal processing.
On the contrary, in D (stability stage), the model explained that “The logs indicate that the ZooKeeper server is actively managing client connections and sessions. It handles new connection attempts, establishes and terminates sessions, and processes session expirations. The server also deals with connection issues, such as end-of-stream exceptions and connection requests from outdated clients.”. This is in line with our analysis on the previous section. Note that the end-of-stream exceptions do not result in server crashes. Further investigation about the seriousness of the end-of-stream exceptions result in the model explaining that they are relatively common and not problematic, as opposed to the checkpoint in C.
@@ IV-C2 Pairwise point comparison
The objective of pairwise comparison is to describe the differences (or similarities) between two given checkpoints. This is helpful when comparing the behavior of the system over time or to correlate visual differences in the Time Curve with the actual set of logs.
First, we selected two close checkpoints from D (stable processing) to find out if the difference is significant. The model explains that both checkpoints are nearly identical except for the presence of a single additional log record related to a KeeperException where an attempt to create an already existing node was made. When further inquiring, the model describes the exception as a non-critical error which could be handled more gracefully. This is in agreement with our manual investigation, where we found out that no server crashes occur during the high-stability stage (Label D).
Second, we selected C and E from the unstable procession stages. On one hand, regarding the similarities, the model explains that both are “… capturing various operational events and issues related to leader election, session management, and client connections”. On the contrary, regarding the differences, first the model returns a generic (yet true) summarization, namely “the first set of logs includes more detailed information about the server environment and configuration, while the second set includes more details about the election process and issues encountered during synchronization and session management.” along with a list of the differences depending on their severity level (i.e., variations by INFO, WARN, etc.). When a more specific explanation for the set of WARN logs is requested , the model replies “These logs highlight issues related to leader election and synchronization, such as unexpected zxid values and exceptions when following the leader, as well as a specific session closure due to the server not running”, which is in line with our manual analysis where we found out that the Zookeeper service is not running temporarily.